# Adaptive Crowdsourcing Via Self-Supervised Learning

### 24 Jan 2024 v1

@@ D.1 Data-Generating Process Satisfies Problem-Formulation Assumptions
D.2 Hyperparameter Tuning for Just-Predict-Others
We describe how we do hyperparameter tuning for this just-predict-others here. For each $K$, we do a separate hyperparameter search. We tune the hyperparameters for the SSL step, and then we tune the hyperparameter for the aggregation step using the best hyperparameters from the SSL step.
@@ D.2 Hyperparameter Tuning for Just-Predict-Others
To tune hyperparameters for both EM and just-predict-others, we assume access to the clairvoyant distribution $\mathbb{P}(Y_{t}\in\cdot|\theta)$. We do this for simplicity. In Appendix D.5, we discuss how hyperparameter tuning can be performed without access to $\mathbb{P}(Y_{t}\in\cdot|\theta)$.
@@ D.3 Estimating MSE
  • To perform hyperparameter tuning for EM and just-predict-others, we sample sixty values of $\Sigma_{*}$.

### 02 Feb 2024 v2

@@ D.1 Data-Generating Process Satisfies Problem-Formulation Assumptions
D.2 Hyperparameter Tuning for Predict-Each-Worker
We describe how we do hyperparameter tuning for this predict-each-worker here. For each $K$, we do a separate hyperparameter search. We tune the hyperparameters for the SSL step, and then we tune the hyperparameter for the aggregation step using the best hyperparameters from the SSL step.
@@ D.2 Hyperparameter Tuning for Predict-Each-Worker
To tune hyperparameters for both EM and predict-each-worker, we assume access to the clairvoyant distribution $\mathbb{P}(Y_{t}\in\cdot|\theta)$. We do this for simplicity. In Appendix D.5, we discuss how hyperparameter tuning can be performed without access to $\mathbb{P}(Y_{t}\in\cdot|\theta)$.
@@ D.3 Estimating MSE
  • To perform hyperparameter tuning for EM and predict-each-worker, we sample sixty values of $\Sigma_{*}$.
@@ D.5 Hyperparameter Tuning with Unobserved Outcomes
Appendix E Extension of Predict-Each-Worker for Settings with Observed Outcomes
Here, we describe an extension for our algorithm that can be used when some or all outcomes are observed. However, if a large fraction of outcomes are observed, then one might be better off using standard supervised learning methods to learn a mapping from estimates to an outcome.
In our extension, we keep the SSL step the same, but we change the aggregation step. First, we train the SSL models on all estimate vectors. Then, for all estimate vectors whose outcomes are known, and for each SSL model, we compute the following statistics: expected error and SSL gradient – this is similar to what we do in Section 3.1. These SSL statistics and the corresponding outcomes are then used to train a neural network. This neural network takes a set of expected errors and SSL gradients as input and produces an estimate of the true outcome. The rationale for using these SSL statistics instead of the estimate vectors as inputs is that it allows leveraging all observed estimates, even the ones for which the outcomes are not observed. We will denote the neural network obtained after training as $\mathfrak{f}$.
Now, we discuss how to produce a group estimate for a new estimate vector $Y_{t}$. We first compute a set of expected errors and SSL gradients based on $Y_{t}$. Then, we input these SSL statistics to neural network $\mathfrak{f}$ to produce an estimate $\hat{Z}_{t}^{(\mathfrak{f})}\in\Re$ of the outcome. Additionally, we use these statistics to compute $\hat{\nu}_{t}^{\top}Y_{t}$, where $\hat{\nu}_{t}$ is based on our aggregation formula (2). Finally, we compute the group estimate by taking a convex combination of $\hat{Z}_{t}^{(\mathfrak{f})}$ and $\hat{\nu}_{t}^{\top}Y_{t}$,
$\hat{Z}_{t}\leftarrow\frac{a}{a+t}\hat{Z}_{t}^{(\mathfrak{f})}+\left(1-\frac{a}{a+t}\right)\hat{\nu}_{t}^{\top}Y_{t}.$
Here, $a\in\Re_{+}$ is a hyperparameter that can be tuned based on a validation set. We should expect $a$ to increase with $t_{o}$, the number of estimates for which the outcome is known. Reason being that larger the value of $t_{o}$, the more we can trust $\hat{Z}_{t}^{(\mathfrak{f})}$